How many grams of water are formed from 20 grams of C5H12 by using this equation C5H12+8O and 5CO2+6H2O
1
Expert's answer
2017-02-16T08:51:12-0500
C5H12+8O = 5CO2+6H2O

n = m/M m = M·n

n (H2O) = 6·n (C5H12)

M (C5H12) = 72 g/mol

M (H2O) = 18 g/mol

n (C5H12) = 20/72 = 0.28 g/mol

n (H2O) = 6·n (C5H12) = 6 · 0.28 = 1.67 g/mol

m (H2O) = 1.67 · 18 = 30 g